OpenStudy (jkristia):

I'm stuck tryingto find a derivate of a pretty simple expression. I will post the expression next

OpenStudy (jkristia):

\[y = \frac{9}{x^2+9}\] Find y' and y'', then find the inflection points I have y' = \[y' = \frac{-18x}{(x^2+9)^2}\] but I'm having trouble getting y''. I have tried using both the quetient rule and the product rule, and in both cases I end up with the same \[y'' = \frac{54(x^2-3)}{(x^2+9)^3}\] argh... found my problem while writing this. I had 54(x^2+3) in the numerator, and the correct answer for the inflection points i +- sqrt(3). I found the problem and now I do have (x^2 - 3), so I get the points to sqrt(3).. Some times it helps to ask the questionin order to find your own mistake